OpenStudy (anonymous):

Howmany atoms are contained in 42.03 grams of Carbon? (scientific notation with 4 significant digits) Please Help!!!!

OpenStudy (mos1635):

N=n*Na

OpenStudy (anonymous):

What?

thomaster (thomaster):

calculate how many moles 42.03 grams of carbon is To do this you divide 42.03 by the molecular mass of carbon (you can find it on http://ptable.com/ ) then multiply it with avogadro's number (6.022*10^23)

OpenStudy (anonymous):

So it should be 42.03 g C * 1/12.011 *( 6.022*10^23)?

thomaster (thomaster):

yes \(\sf\Large\frac{42.03 g~C}{12.011}*\large(6.022*10^{23})\)

OpenStudy (anonymous):

i did the calculation but i got the wrong answer , i got 2.107 *10 ^23 , the answer sheet showed me otherwise which is 1.806*10^24 I'm lost

thomaster (thomaster):

I'll recalculate it: 42.03 gram Carbon MW carbon = 12.011 So \(\sf\large\frac{42.03}{12.011}\)= 3.499 mol carbon From mol to number of atoms = amount of moles multiplied by Na (avogatro's number) \(\sf3.499 * 6.022*10^{23}\) = \(\underline{\color{red}{\sf 2.107*10^{24}}}\) carbon atoms. Somethings wrong with your answer sheet.

OpenStudy (anonymous):

Ok thank you.